# Edge subdivisions and the $L^{2}$-homology of right-angled Coxeter groups

### v1

@@ 1 Introduction
We now explain why computing the $L^{2}$-homology of $\Sigma_{bL}$ for barycentric subdivisions is easier than for general flag triangulations. The program developed in [do01] relied on Mayer–Vietoris arguments, where a flag triangulation $L$ is cut into pieces along links of vertices: $L=L-v\cup_{\operatorname{Lk}v}\operatorname{St}v$. This leads to a decomposition of $\Sigma_{L}$ into pieces whose components are copies of $\Sigma_{L-v}$ and $\Sigma_{\operatorname{St}v}$ intersecting along copies of $\Sigma_{\operatorname{Lk}v}$. For barycentric subdivisions, the link structure of vertices is well known: they decompose as joins of barycentric subdivisions of lower dimensional complexes. Therefore, via Künneth formulas in $L^{2}$-homology they often can be shown to have trivial $L^{2}$-homology, hence by Mayer–Vietoris we can remove them from $L$ without changing $b_{\ast}^{(2)}(\Sigma_{L})$. For barycentric subdivisions of $S^{2n-1}$, for $n\leq 4$, there are enough vertices with $\Sigma_{\operatorname{Lk}(v)}$ $L^{2}$-acyclic such that the resulting complex has dimension $n-1$. Therefore, the Davis manifolds have the same $L^{2}$-Betti numbers as an $n$-dimensional subcomplex; so in particular the $L^{2}$-homology of $\Sigma_{L}$ vanishes above the middle dimension and Poincaré duality implies it vanishes below the middle dimension.
@@ 3.2 Mayer–Vietoris arguments
$\dots\to H_{i+1}^{C_{L}}(W_{L}\times_{W_{\operatorname{Lk}v}}\Sigma_{\operatorname{Lk}v})\to H_{i}^{C_{L}}(W_{L}\times_{W_{\operatorname{St}v}}\Sigma_{\operatorname{St}(v)})\oplus H_{i}^{C_{L}}(W_{L}\times_{W_{L-v}}\Sigma_{L-v})\to H_{i}^{C_{L}}(\Sigma_{L})\to\dots$
@@ 4 ${\mathbb{F}}$-$L^{2}$-homology of RACG’s based on subdivisions
We now use the decompositions of $\Sigma_{L}$ in Lemma 3.2 and Theorem 3.3 to compute $b_{\ast}^{(2)}(W_{L},{\mathbb{F}})$ for certain flag complexes $L$ arising from subdivision procedures.
@@ 4.3 Relative barycentric subdivisions via edge subdivisions
More general versions of Lemmas 4.11 and 4.12 and Corollary 4.14 follow from earlier work of Feichtner–Müller [fm04], Postnikov–Reiner–Williams [pnw08], and Volodin [v10]. Given a simplicial complex $L$, a building set is a collection of simplices $\mathcal{B}$ containing $L^{(0)}$ such that if $\sigma,\tau\in\mathcal{B}$ with $\sigma\cup\tau\in L$ then $\sigma\cup\tau\in\mathcal{B}$ if $\sigma\cap\tau\neq\emptyset$. Any building set determines a corresponding nested set complex, which ends up being a subdivision of $L$ where new vertices correspond to elements in $\mathcal{B}\setminus L^{(0)}$. If $K\subset L$ is a subcomplex, it is easy to see that the vertices of $L$ together with all simplices not contained in $K$ forms a building set, and the corresponding nested set complex is isomorphic to $b(L,K)$.
@@ 5.1 Universal coefficients and torsion growth
We will also let $t_{i}(G)_{p}:=\dim_{{\mathbb{F}}_{p}}\operatorname{Tor}(H_{i}(G;{\mathbb{Z}}),{\mathbb{F}_{p}})$ denote the number of ${\mathbb{Z}}_{p^{k}}$-summands in $H_{i}(G;{\mathbb{Z}})$, and set
@@ 6 Minimally branching complexes
Suppose that $L$ is an $(n-1)$-dimensional flag complex such that every $(n-2)$-simplex is contained in at most three $(n-1)$-simplices. Then $b_{n}^{(2)}(W_{L};{\mathbb{F}})=0$ unless $L$ is an $n$-fold join of $3$ points.
@@
Let $3^{*n}$ denote the $n$-fold join of $3$ points. The theorem obviously holds if $L$ is 00-dimensional, so let $L$ be an $n$-dimensional flag complex with $n\geq 1$ which satisfies the hypotheses, $L\neq 3^{*n+1}$, and suppose that the statement holds for all such $(n-1)$-dimensional flag complexes. We can furthermore assume that $L$ is connected; otherwise $W_{L}$ is a free product of its components. We can also assume that link of each vertex in $L$ is $3^{*n}$; otherwise by induction it would have $b_{n}^{(2)}(W_{\operatorname{Lk}(v)};{\mathbb{F}})=0$ and hence its removal does not affect $b_{n+1}^{(2)}(W_{L};{\mathbb{F}})$ by Lemma 3.2.
Consider one of the $3$-points factors of the link of some vertex. These three vertices are pairwise non-adjacent, and their links cannot be the same copy of $3^{*n}$, since $L$ is connected and $L\neq 3^{*(n+1)}$. Thus we obtain two non-adjacent vertices $v$ and $w$ with $\operatorname{Lk}_{L}v\neq\operatorname{Lk}_{L}w$, and since each link is $3^{*n}$, this implies that $\operatorname{Lk}_{L}w\smallsetminus\operatorname{Lk}_{L}v$ and $\operatorname{Lk}_{L}v\smallsetminus\operatorname{Lk}_{L}w$ are non-empty,
@@
Now, $L$ is a subcomplex of the suspension $S(L-v-w)$ which has $b_{\ast}^{(2)}(W_{S(L-v-w)},{\mathbb{F}})=0$ (we think of $v$ and $w$ as the two suspension points). By repeated application of Theorem 3.3, we will be done if we can iteratively remove the extra edges of $S(L-v-w)$ until we get to $L$, such that at each step, $b_{n}^{(2)}(W_{\operatorname{Lk}e};{\mathbb{F}})=0$.
@@
We first remove edges containing $v$, starting with an edge connecting $v$ to any vertex $u\in\operatorname{Lk}_{L}w\smallsetminus\operatorname{Lk}_{L}v$. The link of this edge is the same as the link of $u$ in $L-v-w$, so it is $3^{*n}$ with a vertex removed, and hence $b_{n}^{(2)}(W_{\operatorname{Lk}e};{\mathbb{F}})=0$ by induction. We continue removing all edges emanating from $v$ and ending outside of $\operatorname{Lk}_{L}v$. Since $L$ is connected, any proper subcomplex of $L$ contains a vertex whose link is a proper subcomplex of $3^{*n}$, so we can iteratively do this for all the extra edges emanating out of $v$.
This does not change any of the links of edges containing $w$, so we do the same procedure starting at a vertex in $\operatorname{Lk}_{L}v\smallsetminus\operatorname{Lk}_{L}w$, until we are left with $L$. ∎

### v2

@@ 1 Introduction
We now explain why computing the $L^{2}$-homology of $\Sigma_{bL}$ for barycentric subdivisions is easier than for general flag triangulations. The program developed in [do01] relied on Mayer–Vietoris arguments, where a flag triangulation $L$ is cut into pieces along links of vertices: $L=L-v\cup_{\operatorname{Lk}v}\operatorname{St}v$. This leads to a decomposition of $\Sigma_{L}$ into pieces whose components are copies of $\Sigma_{L-v}$ and $\Sigma_{\operatorname{St}v}$ intersecting along copies of $\Sigma_{\operatorname{Lk}v}$. For barycentric subdivisions, the link structure of vertices is well known: they decompose as joins of barycentric subdivisions of lower dimensional complexes. Therefore, via Künneth formulas in $L^{2}$-homology they often can be shown to have trivial $L^{2}$-homology, hence by Mayer–Vietoris we can remove them from $L$ without changing $b_{\ast}^{(2)}(\Sigma_{L})$. For barycentric subdivisions of $S^{2n-1}$, for $n\leq 4$, there are enough vertices with $\Sigma_{\operatorname{Lk}v}$ $L^{2}$-acyclic such that the resulting complex has dimension $n-1$. Therefore, the Davis manifolds have the same $L^{2}$-Betti numbers as an $n$-dimensional subcomplex; so in particular the $L^{2}$-homology of $\Sigma_{L}$ vanishes above the middle dimension and Poincaré duality implies it vanishes below the middle dimension.
@@ 3.2 Mayer–Vietoris arguments
$\dots\to H_{i+1}^{C_{L}}(W_{L}\times_{W_{\operatorname{Lk}v}}\Sigma_{\operatorname{Lk}v})\to H_{i}^{C_{L}}(W_{L}\times_{W_{\operatorname{St}v}}\Sigma_{\operatorname{St}v})\oplus H_{i}^{C_{L}}(W_{L}\times_{W_{L-v}}\Sigma_{L-v})\to H_{i}^{C_{L}}(\Sigma_{L})\to\dots$
@@ 4 ${\mathbb{F}}$-$L^{2}$-homology of RACG’s based on subdivisions
We now use the decompositions of $\Sigma_{L}$ in Lemma 3.2 and Theorem 3.3 to compute $b_{\ast}^{(2)}(W_{L};{\mathbb{F}})$ for certain flag complexes $L$ arising from subdivision procedures.
@@ 4.3 Relative barycentric subdivisions via edge subdivisions
More general versions of Lemmas 4.11 and 4.12 and Corollary 4.14 follow from earlier work of Feichtner–Müller [fm04], Postnikov–Reiner–Williams [pnw08], and Volodin [v10]. Given a simplicial complex $L$, a building set is a collection of simplices $\mathcal{B}$ containing $L^{(0)}$ such that if $\sigma,\tau\in\mathcal{B}$ with $\sigma\cup\tau\in L$ then $\sigma\cup\tau\in\mathcal{B}$ if $\sigma\cap\tau\neq\emptyset$. Any building set determines a corresponding nested set complex, which ends up being a subdivision of $L$ where new vertices correspond to elements in $\mathcal{B}\smallsetminus L^{(0)}$. If $K\subset L$ is a subcomplex, it is easy to see that the vertices of $L$ together with all simplices not contained in $K$ forms a building set, and the corresponding nested set complex is isomorphic to $b(L,K)$.
@@ 5.1 Universal coefficients and torsion growth
We will also let $t_{i}(G)_{p}:=\dim_{{\mathbb{F}}_{p}}\operatorname{Tor}(H_{i}(G;{\mathbb{Z}});{\mathbb{F}_{p}})$ denote the number of ${\mathbb{Z}}_{p^{k}}$-summands in $H_{i}(G;{\mathbb{Z}})$, and set
@@ 6 Minimally branching complexes
If $L$ is a connected flag graph with vertices of degree $\leq 3$, then $b_{2}^{(2)}(W_{L};{\mathbb{F}})=0$ unless $L=K_{3,3}$.
@@
If $L$ has a vertex of degree $\leq 2$, we can iteratively remove the vertices of $L$, at each stage removing a vertex of degree $\leq 2$. These removals does not affect $b_{2}^{(2)}(;{\mathbb{F}})$, hence $b_{2}^{(2)}(W_{L};{\mathbb{F}})=0$. Therefore, we can assume that $L$ is trivalent. Furthermore, note that if $L$ has a proper full subcomplex $A$ so that $L-A$ is disconnected and $b_{1}^{(2)}(W_{A};{\mathbb{F}})=0$, then we get a decomposition $L=L_{1}\cup_{A}L_{2}$ where each $L_{i}$ has a vertex of degree $\leq 2$. Therefore, $b_{2}^{(2)}(W_{L_{i}};{\mathbb{F}})=0$ so $b_{2}^{(2)}(W_{L};{\mathbb{F}})=0$ by Mayer–Vietoris. In particular, we are done if $L$ has a separating edge or separating pair of non-adjacent vertices, so assume that this does not occur and that $L\neq K_{3,3}$.
We now show that $L$ has two non-adjacent vertices $u$ and $v$ with non-separating stars. There are two cases: either $L$ contains a copy of a suspension of three points as a full subcomplex, or it does not.
Suppose there is a suspension of three points inside of $L$ with $x$ and $y$ suspension points and base $\{a,b,c\}$. Let $\{a^{\prime},b^{\prime},c^{\prime}\}$ denote the endpoints of edges starting at $\{a,b,c\}$ which are not in the suspension. Note that $\{a^{\prime},b^{\prime},c^{\prime}\}$ are all different. Indeed, since $L\neq K_{3,3}$, they cannot be all the same, and if $a^{\prime}=b^{\prime}$, then $a^{\prime}$ and $c^{\prime}$ would separate the suspension from the rest of the complex.
Next we claim that the stars of $\{a,b,c\}$ are all non-separating. Consider a vertex $z\in L-\operatorname{St}a-c$. Since the pair $a^{\prime},c$ does not separate, there is a geodesic in $L-a^{\prime}-c$ from $z$ to $b$. Such geodesic cannot go through $x$, $y$ and $a$, so $z$ is connected to $b$ in $L-\operatorname{St}a-c$. Thus $L-\operatorname{St}a-c$ is connected, and since $c$ is adjacent to $c^{\prime}$, $L-\operatorname{St}a$ is connected as well. Therefore we can take $a$ and $b$ as $u$ and $v$.
For the second case, we claim that we can take $u$ and $v$ to be points of maximum distance. The maximality of the distance implies that for any vertex in $\operatorname{Lk}v$ at least one of the edges from it points in the direction of $u$. Suppose, to the contrary, that $\operatorname{St}v$ separates, and let $C$ be a component of $L-\operatorname{St}v$ not containing $u$. Then the first edge of a geodesic in $L$ from any vertex of $C$ to $u$ must have the endpoint in $\operatorname{Lk}v$. Therefore, vertices in $\operatorname{Lk}v$ are adjacent to at most one vertex in $C$, and vertices in $C$ are all adjacent to a vertex in $\operatorname{Lk}v$. So, $C$ has at most three vertices. $C$ cannot be a single vertex, since $L$ does not contain suspensions of three points. If there are two vertices in $C$, then they are adjacent and each is adjacent to two vertices in $\operatorname{Lk}v$, so one vertex in the link is adjacent to both, contradiction. If there are three, then each is adjacent to a unique vertex in the link, since $L$ is trivalent they all must be adjacent, so we get a triangle, contradicting flagness. Thus stars of $u$ and $v$ do not separate, and we are done with the second case.
@@
Now, $L$ is a subcomplex of the suspension $S(L-u-v)$ which has $b_{\ast}^{(2)}(W_{S(L-u-v)};{\mathbb{F}})=0$ (we think of $u$ and $v$ as the two suspension points). By repeated application of Theorem 3.3, we will be done if we can iteratively remove the extra edges of $S(L-u-v)$ until we get to $L$, so that at each step, $b_{1}^{(2)}(W_{\operatorname{Lk}e};{\mathbb{F}})=0$.
@@
We first remove edges containing $u$ which join $u$ to vertices $\notin\operatorname{Lk}_{L}u$, starting with an edge whose other endpoint is $w\in\operatorname{Lk}_{L}v\smallsetminus\operatorname{Lk}_{L}u$. (Note that since $u$ and $v$ are note adjacent and $L\neq K_{3,3}$ the links of $u$ and $v$ are different.) The link of this edge is the same as the link of $w$ in $L-u-v$, so it is $2$ points and hence $b_{\ast}^{(2)}(W_{\operatorname{Lk}e};{\mathbb{F}})=0$. We continue removing all edges emanating from $u$ and ending outside of $\operatorname{Lk}_{L}u$. Since $L-\operatorname{St}u$ is connected, any vertex in $L-\operatorname{St}u$ is connected by a path to $v$, and hence connected to a vertex in $\operatorname{Lk}v$. Therefore, we can always remove these edges so that links are $\leq 2$ points. Removing these edges containing $u$ does not change any of the links of edges containing $v$, so we do the same procedure starting at a vertex in $\operatorname{Lk}_{L}u\smallsetminus\operatorname{Lk}_{L}v$, until we are left with $L$. ∎
The higher dimensional version of Theorem 6.1 is easier to prove, due to the following lemma. Let $3^{*n}$ denote the $n$-fold join of $3$ points with itself.
Lemma 6.2.
Suppose $L$ is an $n$-dimensional connected flag complex with links of vertices $\cong 3^{*n}$. If $n\geq 2$, then $L=3^{*n+1}$.
Proof.
Let $v\in L$, and let $x\in\operatorname{Lk}v$. Then, since $\operatorname{Lk}(x)\cong 3^{*n}$, $x$ is adjacent to two vertices $u$ and $w$ that are not in $\operatorname{St}v$, and $\{v,u,w\}$ is a $3$ point factor of $\operatorname{Lk}(x)$. Hence, if $y\in\operatorname{Lk}(v)$ is adjacent to $x$, both $u$ and $w$ are adjacent to $y$. Since $\operatorname{Lk}v\cong 3^{*n}$ is connected for $n\geq 2$, it follows that $u$ and $w$ are the same for all vertices in $\operatorname{Lk}v$. So we get a copy of $3^{*n+1}$ in $L$, and since $L$ is connected we are done. ∎
Theorem 6.3.
Suppose that $L$ is an $(n-1)$-dimensional flag complex such that every $(n-2)$-simplex is contained in at most three $(n-1)$-simplices and links of $k$-simplices are connected for $-1\leq k<n-2$. Then $b_{n}^{(2)}(W_{L};{\mathbb{F}})=0$ unless $L=3^{*n}$.
Proof.
The proof is by induction on $\dim L$. If $L$ is 00-dimensional it is obvious and if $L$ is $1$-dimensional, then it is Theorem 6.1. Now assume that the statement holds for all such $(n-1)$-dimensional flag complexes, and let $L\neq 3^{*n+1}$ be an $n$-dimensional flag complex with $n\geq 1$ which satisfies the hypotheses. Note that $L$ is connected and if $v$ is a vertex of $L$ then $\operatorname{Lk}v$ satisfies the conditions of the theorem, since the link of $\sigma$ in $\operatorname{Lk}v$ is the link of $v\sigma$ in $L$.
We can assume that the link of each vertex in $L$ is $3^{*n}$. Otherwise by induction it would have $b_{n}^{(2)}(W_{\operatorname{Lk}v};{\mathbb{F}})=0$, and hence its removal does not affect $b_{n+1}^{(2)}(W_{L};{\mathbb{F}})$ by Lemma 3.2. So, we are done by Lemma 6.2. ∎